A certain unfair coin lands on heads 1/3 of the time. The coin is tossed 5 times. Finds the probability of no more than 3 heads


Answer by swincher4391(1107)

Probability of heads = 1/3
Probability of tails = 2/3
P[no more than 3 heads] = 1-P[4 heads or 5 heads] = 1-(P[4 heads] + P[5 heads])
1 - ((5 choose 4)(1/3)^4*(2/3)^1 + (5 choose 5)(1/3)^5(2/3)^0) = 0.9547